OpenStudy (heart_offire):

Factor the polynomial. x^2 - 6x + 9 A) (x - 9)(x - 1) B) (x - 3)(x + 3) C) (x + 9)(x + 1) D) (x - 3)(x - 3)

OpenStudy (heart_offire):

i think its B but.....

OpenStudy (mathmale):

Why do you think it's B? If you take the two factors listed in B, and multiply them together, what do y ou get?

OpenStudy (heart_offire):

i get positive 9

OpenStudy (heart_offire):

and there's a positive 9 in the equation so i thought it was right

OpenStudy (mathmale):

You've multiplied 3*3. No. Instead, multiply the possible factors together: (x - 3)(x + 3) = ?

OpenStudy (heart_offire):

well the possible factors for 3 in addition is 2,1

OpenStudy (heart_offire):

but if i multiply 2 and 1 i'll 2

OpenStudy (mathmale):

Heart: I need for you to multiply. You're factoring. No. Please multiply: (x+3)(x-3)=?

OpenStudy (heart_offire):

(x^2+9)?

OpenStudy (mathmale):

Use the "FOIL" method. (x+3)(x-3). x*x=? 3*x+? x*3=? (-3)(3)=?

OpenStudy (heart_offire):

x*x=x^2 3*x=3x x*3=x3 (-3)(3)=-9

OpenStudy (mathmale):

You are multiplying two binomials together: (x-3)(x+3)=x*x-3x+3x-9. Please simplify this.

OpenStudy (mathmale):

OpenStudy says you're focusing on some other problem. I'm sorry our conversation has gone on and on, but I believe all of it has been important and necessary. Do you want to finish this problem, or drop it and go to the other one?

OpenStudy (heart_offire):

i honestly dont know how to solve this problem and i have other problems posted but im not occupied by them now. i want to finish this problem

OpenStudy (mathmale):

You originally chose B as your answer, so I have been leading you through the multiplication of (x+3)(x-3) so that you can determine whether B is actually correct or not. It is not. Why? because the product is x^3 - 9, which is not the same as x^2 -6x + 9.

OpenStudy (mathmale):

So you must eliminate choice B. Which other choice might you pick? How would you check whether or not the factors given in that choice multiply out to x^2 - 6x + 9?

OpenStudy (mathmale):

Hint: Multiply: (x-3)(x-3)

OpenStudy (mathmale):

Perhaps you should take a break and then return to solving this problem. I need to take one myself. Being able to multiply binomials, such as to multiply (x-3)(x-3), is a very important skill that you will need in many applications of algebra.

OpenStudy (heart_offire):

sure

OpenStudy (heart_offire):

OHHHH i think i get now......x^2-3x+9....is that the answer

OpenStudy (heart_offire):

so are we going to finish the problem

OpenStudy (carlyxphan):

The answer is D (See picture for explanation)

OpenStudy (susanna):

x^2-6x+9 (x )(x ) (x - )(x - ) Both negative because 9 is positive and -6 is negative, and there 2 numbers should add up to -6, and when multiplied, should be 9. (x - 3)(x - 3) FOIL to check First times first: x*x x^2 Outsides multiplied: x*-3 x^2-3x Insides multiplied: -3*x x^2-3x-3x Last times last: -3*-3 x^2-3x-3x+9 Now, add like terms: \[x ^{2}-6x+9\]

OpenStudy (heart_offire):

thanks that makes sense